# Removing the need for ground truth UWB data collection: self-supervised ranging error correction using deep reinforcement learning

### v1

@@ VI-B Training
The RL algorithm was trained for 1000 episodes with $\gamma~{}=~{}0.5$, $\tau_{critic}~{}=\tau_{actor}~{}=~{}0.01$, $\alpha_{actor}~{}=~{}5e^{-5}$ and $\alpha_{critic}~{}=~{}5e^{-4}$. The patience of the learning rate schedulers was set to 150 episodes.
@@
Figure 6: Trajectory comparison of the original smoothed KF (with no RL correction) with the improved trajectories during training after 400 and 1000 episodes
Figure 6 illustrates the iterative improvement by the RL algorithm during training. The green curve represents the smoothed KF trajectory without any RL correction, this is the data used to calculate the reward before the target actor is updated. Once the target actor starts getting updates from episode 350 onwards, the data used to calculate the rewards starts improving. The orange trajectory is used for reward calculation at episode 400 and the green trajectory is the improved trajectory at episode 1000, the end of training. This visually illustrates the iterative improvement of the algorithm.
@@ VI-C Evaluation
In Figure 7, the box plots show that the proposed RL algorithm performs better than the supervised CNN approach and significantly better than the uncorrected UWB. The median range error of the RL algorithm is lower than the other two methods for both box plots.
A first explanation is that our approach is rewarded when the overall trajectory improves. As such, our approach focuses on those CIR types that most negatively impact the overall trajectory. In contrast, traditional CNNs optimize all CIRs equally, regardless of their importance to the overall localization. A second explanation relates to an improved design of the (actor) neural network compared to prior work. Our last layer has Tanh activation with output scaling while the prior CNN uses no activation in the last layer so that the output could be any real number which might lead to numerical instability in some cases.
Moreover, the ability to learn online from new experiences without supervision means the RL agent can seamlessly adapt to changes in environments by continuing the training process, shown in Section VI-D, the RL approach can learn the specifics of an environment even if the environment changes. Additionally, the range errors of the RL algorithm are more tightly clustered around the median. This indicates that the RL algorithm is more robust in its performance than the other two methods. Figure 7(b) highlights the performance of the proposed RL algorithm in NLOS situations, with a lower median error rate and smaller interquartile range, this again indicates a more consistent performance in the more difficult situations. However, it has visibly more outliers than the supervised CNN approach. Table IV tells a similar story. The proposed RL approach reduces MAE by 31.6% compared to uncorrected UWB and by 14.5% compared to the supervised CNN. The separate NLOS results emphasize the increased performance even more, as the proposed algorithm decreases the MAE by 34.8% compared to uncorrected UWB and by 22.8% compared to the supervised CNN.
TABLE IV: Quantitative results of the baselines and proposed algorithm
Method MAE (mm) all samples MAE (mm) NLOS
Uncorrected UWB 155 181
@@
Proposed RL algorithm 106 118
@@
Figure 7: The ranging errors of uncorrected UWB, the supervised CNN and our proposed RL algorithm during evaluation for (a) all samples and (b) only in NLOS samples. The figures show that our proposed self-supervised RL algorithm performs comparable or better than a supervised CNN approach.
VI-D Adaptivity of algorithm
@@
Figure 8: MAE comparison of our approach with a state-of-the-art CNN for error correction [] for a changing environment. The proposed RL algorithm is compared with uncorrected UWB and the supervised CNN trained on the current environment and the other environment. The figure shows the generalization problem of the supervised CNN and that the proposed RL algorithm can adapt itself to a changing environment.
@@
The same environment can change over time. To evaluate the performance of our algorithm when there are sudden changes in the environment, a new dataset was collected in the same warehouse 6 months later. At that time, there were more goods in the racks, additional clutter in the warehouse (obstacles, boxes, …) and the anchor nodes experienced many small disturbances over time. These combined effects lead to a more challenging environment, which is reflected in the higher MAE for uncorrected UWB. The MAE during training is shown in Figure 8. First, the proposed algorithm is trained on ”environment 1”, which is the original dataset used in the previous evaluation, between episodes 0-500, the training is similar to Figure 5, except it is halted after 500 instead of 1000 episodes. After 500 episodes, the environment is switched to ”environment 2”, the same environment but 6 months later and more difficult as discussed before. The learning rates of the RL algorithm are reset to the starting values and the exploration is increased. At first, the RL algorithm leads to worse performance, due to the exploration, but quickly adapts to the environment and surpasses the supervised model trained on the first dataset and later also the supervised CNN trained on the new dataset. This result displays the adaptivity of the RL algorithm compared to the supervised CNN approach, without needing to label a dataset it can adapt to changing environments and continuously lead to better ranging performance. The supervised CNN approach requires a new labeled dataset in a changed environment, while our approach does not.
VI-E Complexity analysis
VI-E1 Algorithmic complexity
@@
VI-E2 Time complexity
The algorithm’s convergence time duration is limited by the incoming data rate of the UWB localization system. The UWB localization has a sampling rate of 50 Hz, because the dataset has about 3000 samples the time duration of 1 episode is about one minute.
@@ VII Future work
There are several avenues to further expand on this research. A first potential enhancement could be to make the Kalman Filter adaptive to the trajectory. Straight trajectories can have more smoothening, while corners need reduced smoothening. By modifying the smoothening and filtering process based on trajectory characteristics, the tracking system’s accuracy and robustness can be further investigated. A second area of improvement lies in the selection of points from the smoothed and filtered trajectory to the discrete data points for learning. Currently, the middle point in the buffer is associated with the average position. Future research could investigate the feasibility of defining a continuous trajectory and selecting the closest point. This adjustment could potentially lead to more responsive and accurate error correction processes. Additionally, the system’s capabilities could be expanded by integrating various sources of additional information. This includes exploring adding map data, reflections, CIR, and range data between anchor nodes (with known fixed positions). Furthermore, Inertial Measurement Unit (IMU) data could be added as input to the Kalman Filter to make it more robust and allow for better labels. Finally, applying this research to positioning systems using TDoA methods instead of TWR systems. TDoA, which relies on measuring the time delays of signals arriving at different nodes, is a widely used technique in wireless localization, and applying this research there would further improve the real-world practicality of more positioning systems.

### v2

@@ VI-B Training
The RL algorithm was trained for 1000 episodes with $\gamma~{}=~{}0.5$, $\tau_{critic}~{}=\tau_{actor}~{}=~{}0.01$, $\alpha_{actor}~{}=~{}5e^{-5}$ and $\alpha_{critic}~{}=~{}5e^{-4}$. The patience of the learning rate schedulers was set to 150 episodes. The batch size is 50 and the buffer size $N$ is set to 31 as this was shown to be a good trade-off. To evaluate the performance of our machine learning models, we split the dataset into training and testing sets. Specifically, we used an 80/20 split, where 80% of the data was allocated for training the models, and the remaining 20% was reserved for validation. Typically, RL does not require a distinct training and validation split, as the performance is usually evaluated based on the agent’s interaction with the environment. However, since we are working with a predefined dataset and aim to make fair comparisons with supervised learning methods, we implemented this split.
The validation data is still used to maintain consistent trajectories and averaging. However, it is never incorporated into the replay buffer. This means that the actor and critic networks have not trained on the validation CIRs, ensuring that the RL algorithm has not been exposed to the CIR-correction combinations shown in the evaluations. This approach allows us to assess the generalization capabilities of our models. All the following evaluations use the first dataset that is described in section III except in subsection VI-D where both are used.
@@
Figure 6: Trajectory comparison of the original EKF (with no RL correction) with the improved trajectories during training after 300 and 500 episodes
Figure 6 illustrates the iterative improvement by the RL algorithm during training. The green curve represents the smoothed EKF trajectory without any RL correction, this is the data used to calculate the reward before the target actor is updated. Once the target actor starts getting updates from episode 350 onwards, the data used to calculate the rewards starts improving. The orange trajectory is used for reward calculation at episode 300 and the green trajectory is the improved trajectory at episode 500. This visually illustrates the iterative improvement of the algorithm.
@@ VI-C Evaluation
In Figure 7, the box plots show that the proposed RL algorithm performs better than the supervised CNN approach and significantly better than the uncorrected UWB. The median range error of the RL algorithm is lower than the other two methods for the NLOS box plots but slightly higher than the median of the supervised CNN when comparing all evaluation samples. However, the interquartile range (IQR) is lower and thus more tightly clustered around the median. This indicates that the RL algorithm is more robust in its performance than the other two methods. Figure 7(b) highlights the performance of the proposed RL algorithm in NLOS situations, with a lower median error rate and smaller interquartile range, this again indicates a more consistent performance in the more difficult situations. However, it has visibly more outliers than the supervised CNN approach. Table IV tells a similar story. The proposed RL approach reduces MAE by 31.6% compared to uncorrected UWB and equals the supervised CNN from [14]. The separate NLOS results emphasize the increased performance even more, as the proposed algorithm decreases the MAE by 34.8% compared to uncorrected UWB and by 22.8% compared to the supervised CNN. Table IV also contains the results of the optimized supervised CNN that uses the Tanh activation with output scaling instead of the linear layer. These results show that the improved accuracy compared to the supervised approach, visible in figure 5 is mainly due to Tanh activation and scaling.
TABLE IV: Quantitative results of the baselines, the optimized version of the supervised CNN, and the proposed algorithm
Method MAE (mm) MAE (mm)
All Samples NLOS
Uncorrected UWB 162 194
@@
Optimized Supervised CNN 112 137
Proposed RL Algorithm 112 128
@@
Figure 7: The ranging errors of uncorrected UWB, the supervised CNN, and our proposed RL algorithm during evaluation for (a) all samples and (b) only in NLOS samples. The figures show that our proposed self-supervised RL algorithm performs comparable or better than a supervised CNN approach.
VI-D Impact of changing environment
@@
Figure 8: MAE comparison of our approach with a state-of-the-art CNN for error correction [14] for a changing environment. The proposed RL algorithm is compared with uncorrected UWB and the supervised CNN trained on the current environment and the other environment. The figure shows the generalization problem of the supervised CNN and that the proposed RL algorithm can adapt itself to a changing environment.
@@
The same environment can change over time. To evaluate the performance of our algorithm when there are sudden changes in the environment, a new dataset was collected in the same warehouse 6 months later, as discussed in section III. Figure 8 shows the MAE during training. First, the proposed algorithm is trained on ”environment 1”, which is the original dataset used in the previous evaluation, between episodes 0-500, the training is similar to Figure 5, except it is halted after 500 instead of 1000 episodes. After 500 episodes, the environment is switched to ”Environment 2”, which is more difficult. The learning rates of the RL algorithm are reset to the starting values and the exploration is increased. Our reason for resetting the learning rate was to accelerate the learning. The current learning rate was low due to convergence in ”environment 1”. In a new environment, we reset the learning rate because it is not converged anymore. This adjustment can be made during the system’s deployment. When significant changes occur in the environment, such as a new layout or the addition of racks, products, or items, increasing the learning rate allows the system to adapt more rapidly. Conversely, when the environment remains relatively stable, a lower learning rate suffices to accommodate everyday changes. The dotted line in Environment 2 is the accuracy of the supervised CNN trained in Environment 1 but executed in Environment 2. This is a realistic situation in which a trained model is executed in an environment. These results show that traditional CNNs start to degrade over time when normal changes are made to the environment. As such, even using a single environment, we show that adapting the neural network is necessary for real-world usability. Our proposed RL algorithm provides this. At first, the RL algorithm leads to worse performance, due to the exploration, but quickly adapts to the environment and surpasses the supervised model trained on the first dataset and later also the supervised CNN trained on the new dataset. This result displays the adaptivity of the RL algorithm compared to the supervised CNN approach, without needing to label a dataset it can adapt to changing environments and continuously leads to improved ranging performance. The supervised CNN approach requires a new labeled dataset in a changed environment, while our approach does not.
VI-E Impact of robot trajectory
In training our neural network, we assumed the occasional presence of people or vehicles following predictable trajectories. However, after training on these predictable patterns, our neural network can accurately adjust to both individual positions and unpredictable trajectories. This section assesses the model’s performance on a more unpredictable path to demonstrate that its learned behavior generalizes to other movement patterns. The data gathered over this unpredictable path, shown in Figure 9 in black, consists of 300 UWB ranges with a MAE of 173 mm, after correction (using the trained actor model) the MAE is reduced to 128 mm. This improvement is in line with the previously reported improvements showing that our trained model also performs on more random trajectories. This improved ranging performance leads to an improved trajectory visible in Figure 9.
Figure 9: Once trained, our approach can also correct individual positions or unpredictable paths. The figure shows the evaluation of our approach for a trajectory with unpredictable paths, showing the original unpredictable path (in black - ground truth obtained using a MoCAP system), the corrected path using a traditional EKF approach (in blue), and the improved trajectory using our RL correction (in green).
VI-F Complexity analysis
VI-F1 Algorithmic complexity
@@
VI-F2 Time complexity
In terms of time, we have two constraints: (1) time required for (automatic) data capture (2) and time required for training and inference. Time required for (automatic) data capture: we need about 200 episodes to retrain the network from scratch in a new environment (visible in Figure 8). Each episode is about 3000 samples, sampled at 50 Hz, this leads to about 1 minute of data capturing per episode. This means that our approach requires about 3 hours and 20 minutes of data capturing if starting the system from scratch (e.g. after installation). However, while the system is operational, this data is collected automatically, without requiring human intervention. As such, our system can be used in environments where data is collected continuously and keeps adapting at run-time. In contrast, the supervised CNN first necessitates an entire manual measurement campaign to collect a dataset with ground truth, followed by the model training process. The RL approach eliminates the need for a ground truth dataset, encapsulating the entire process within the algorithm itself. The training and inference time was measured on two different hardware platforms. Training one batch containing 50 samples, (both actor and critic network) takes 0.09 seconds on an NVIDIA GeForce GTX 1080 Ti GPU and 0.19 seconds on an Intel(R) Core (TM) i7 8700 CPU. Due to the sampling time of the system being 50Hz, the models can be trained faster than the incoming data rate of the new samples. The inference time of the actor network (providing the corrections) on an NVIDIA GeForce GTX 1080 Ti GPU, is 0.0025 seconds, while on an Intel(R) Core (TM) i7 8700 CPU. Given that the UWB localization system operates at a sampling rate of 50 Hz, a new range and CIR are received every 0.02 seconds. The way our proposed approach works is that each incoming sample first needs a correction from the anchor and then it is trained in batches of samples based on the replay buffer. Generating correction for 50 samples thus takes 0.125 seconds on the GPU and 0.18 seconds on the CPU. Training on this batch takes an additional 0.09 or 0.19 seconds respectively. Combined this leads to a processing time of 0.18 and 0.37 seconds. Processing one second of incoming data takes less than one second. An overview of the timing aspects is given in Table V. In summary, our method offers a more flexible, efficient, and adaptable solution that can operate in real-time and continuously improve its performance without manual intervention, addressing the limitations of traditional CNN approaches in dynamic environments.
Time Aspect Our Method Supervised Method
Initial Data Collection Automatic $\sim$3h 20min Manual campaign (time varies)
Ground Truth Labeling Not required Required (Expensive, time-consuming)
Training Time (per batch) 0.09s (GPU), 0.19s (CPU) Not provided (typically longer)
Inference Time 0.0025s (GPU), 0.0036s (CPU) Not provided
Adaptation to New Environments Continuous, automatic Requires new data collection and training
TABLE V: Comparison of time aspects showing that our method offers a more flexible, efficient, and adaptable solution.
@@ VII Future work
There are several avenues to further expand on this research. A first potential enhancement could be to make the EKF adaptive to the trajectory. Straight trajectories can have more smoothening, while corners need reduced smoothening. By modifying the smoothening and filtering process based on trajectory characteristics, the tracking system’s accuracy and robustness can be further investigated. A second area of improvement lies in the selection of points from the smoothed and filtered trajectory to the discrete data points for learning. Currently, the middle point in the buffer is associated with the average position. Future research could investigate the feasibility of defining a continuous trajectory and selecting the closest point. This adjustment could potentially lead to more responsive and accurate error correction processes. Additionally, the system’s capabilities could be expanded by integrating various sources of additional information. This includes exploring adding map data, reflections, CIR, and range data between anchor nodes (with known fixed positions). Furthermore, Inertial Measurement Unit (IMU) data could be added as input to the EKF to make it more robust and allow for better labels. Finally, apply this research to positioning systems using TDoA methods instead of TWR systems. TDoA, which relies on measuring the time delays of signals arriving at different nodes, is a widely used technique in wireless localization, and applying this research there would further improve the real-world practicality of more positioning systems.